OpenStudy (anonymous):

A car accelerates from rest to a velocity of 5 meters/second in 4 seconds. What is its average acceleration over this period of time?

OpenStudy (loser66):

\[a =\frac{v-v_0}{t}=\frac{5}{4}m/s^2\]